40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 26 20:30:47 2011, Katrin, Update, LSC, POX11 and POP55 installed 
    Reply  Fri Oct 28 03:42:50 2011, kiwamu, Update, LSC, ITMX table needs to be refined ITMXtable.png
Message ID: 5752     Entry time: Fri Oct 28 03:42:50 2011     In reply to: 5743
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: ITMX table needs to be refined 

(POX)

The POX beam had been 80% clipped at a black glass beam dump of the POX11 RFPD.

I steered the first mirror in the POX path to fix the clipping. Then the beam was realigned onto the RFPD.

However the beam is still very close to the black glass, because the incident angle to the second mirror is not 45 deg .

We need to refine the arrangement of the POX11 optics a bit more so that the beam will never be clipped at the black glass.

 

(POP)

 The POP optics also need to be rearranged to accommodate one more RFPD.

Additionally Rana, Suresh and I discussed the possible solutions of POP22/110 and decided to install a usual PD (PDA10A or similar) instead of a custom-made.

So a plan for the POP detectors will be something like this:

        + design an optical layout.

        + buy a 2 inch lens whose focul length is long enough (#5743)

        + rearrange the optics and install POP22/110

        + lay down a long SMA cable which sends the RF signal from POP22/110 to the LSC rack.

        + install a power splitter just before the demod board so that the signal is split into the 22MHz demoad board and 110MHz demod board.

           => make sure we have a right splitter for it.

        + install a band pass filter after the power splitter in each path.

           => A 22MHz band pass filter is already in hand. Do we have 110MHz band pass filter somewhere in the lab ?

The picture here shows the latest configuration on the ITMX table.

ITMXtable.png

Quote from #5743

RF photo diodes POP55 and POX11 are installed. The beams are aligned to the photo diodes. 

ELOG V3.1.3-